OpenStudy (anonymous):

Can some one please help me? will metal and fan t who ever answers..... Help mehhhh I give medals... :3 Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (cwrw238):

let car A travel at x mph speed = distance/time for car A we have x = d / 2.4 so d = 2.4x for car B we have x + 22 = d / 4 so d = 4(x + 22) distances are the same so 4(x + 22) = 2.4x solve this equation for x to get speed for car A.

OpenStudy (cwrw238):

oops - sorry - i made an error speed for car B is x - 22 of course - Car A is the faster

OpenStudy (cwrw238):

so the equation is 4(x - 22) = 2.4x

OpenStudy (cwrw238):

Ok - do you follow that?

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (cwrw238):

sorry gtg right away

OpenStudy (anonymous):

oh